# Attention on flow control: transformer-based reinforcement learning for lift regulation in highly disturbed flows

### v2

@@ III.1 Benchmarking control performance: mid-chord pitching in single-gust environment
We begin by investigating a baseline environment, which is characterized by a single upstream gust. This configuration is set up to be a benchmark for comparing the performance of P control against RL control. Specifically, we intend to evaluate how well each approach can mitigate the lift variation induced by the gust, and to establish a performance baseline before we move on to a more complex environment. The P control strategy, also explored in previous work as a baseline for RL [31], is defined as follows:
@@
We note in passing that, for linear control in this work, we have limited ourselves to proportional control only. It is obviously possible to extend the P control to incorporate additional control terms, such as integral (I) and derivative (D), i.e. PID control. However, for this nonlinear dynamical system the derivation of the control gains requires a complex manual tuning process which cannot be exactly aligned with the reward function structure. In contrast, the RL agent learns the optimal control strategy with respect to the reward function automatically, and this offers a more scalable and adaptive solution that extends to problems with nonlinear dynamics, as we will explore in the following sections.
@@ III.2 Extending control capabilities: mid-chord pitching in multi-gust environment
In Fig. 10 we evaluate the trained RL control strategies against the P control strategies over $45$ random cases with three gusts. First, we observe that the case with no control yields a total reward of around 175-175, which is significantly lower than the corresponding value of approximately $15$ in a single-gust environment and reflects the stronger disturbance introduced by three successive gust. Furthermore, the total reward increases with the $K_{p}$ value until it reaches the peak total reward of around 14-14 at $K_{p}=70$. Beyond this point, the performance starts to deteriorate, likely due to the system becoming overly reactive to the gust. Meanwhile, the RL agent shows a clear improvement in performance from the initial pretrained policy (blue dashed line) to the final pretrained policy (blue envelope). The final RL control surpasses the best P control by a reward margin of approximately $7$. This performance gap is notably larger than the marginal $0.4$ improvement observed in the single-gust environment. This improvement supports the hypothesis that the RL control scales better with increased system complexity and nonlinearity compared to proportional control, and thus demonstrates the advantage of the RL approach in a more challenging multi-gust environment.
@@
In summary, our results confirm that the performance gap between RL and P control widens in the multi-gust environment, validating our hypothesis that RL control is more scalable in handling increased system complexity and nonlinearity. Moreover, we observe that both RL agents trained in single- and three-gust environments can be applied to n-successive gusts effectively, and that the RL agent trained in three gusts generalizes slightly better than the one trained on a single gust, but the margin of improvement is moderate due to the actuation magnitude limit. These insights suggest that further performance gains may be achievable with increased control authority, and thus motivate us to examine the effect of pivot point location on control effectiveness in the next sections.
@@ III.3 Revisiting the pivot location: quarter-chord pitching in single-gust environment
We define the first of these, the added mass contribution, using potential flow theory [48]:
@@
The small $K_{p}$ controller also leads to expert data with minimal action variation, which, when used as pretraining data, offers little advantage over the default (random) initialization of the RL policy network. Specifically, since we use zero-center initialization (a standard PyTorch initialization [49]) for the policy networks, which produces a near-zero output already, and we employ a fixed initial policy standard deviation of $1$, which allows for substantial exploration early in training, the influence of the expert data is diminished during initial policy updates. Therefore, we chose not to use pretraining in this case. To clarify, our decision to omit pretraining in this configuration is not a limitation of the framework, but rather a deliberate design choice driven by the observed ineffectiveness of expert data in this particular setup. As we will demonstrate, even without pretraining, the RL agent is capable of learning an effective control strategy in the quarter-chord configuration. For the readers who may be interested in pursuing pretraining under these conditions, one potential solution is to develop a full PID controller, which would likely be more successful than P control in reducing the overreaction. One could also apply a scaling factor to amplify the action magnitudes from the expert data before using them for pretraining, which would ensure that the pretrained policy deviates more meaningfully from the zero-center initialized baseline. Another approach may be to employ a smaller initial policy standard deviation, which would constrain the exploration range and help maintain the influence of the pretrained expert data. However, these approaches require careful tuning, and are outside the scope of the present study.
@@
Figure 16: The corresponding vorticity contour and surface pressure distribution at the four steps indicated in Fig. 15. $C_{\Delta p}$ is scaled $\frac{1}{4}$ from its original value.
@@
Figure 18: The corresponding vorticity contour and surface pressure distribution at the four steps indicated in Fig. 17. $C_{\Delta p}$ is scaled $\frac{1}{4}$ from its original value.
@@ IV Conclusions
We have found that P control is an effective control strategy in the mid-chord pitching configuration on its own, but that the performance improvement that RL control exhibits over P control widens for multiple-gust scenarios, suggesting that RL control is more scalable in handling system complexity and nonlinearity. In the quarter-chord pitching configuration, we have found that the RL control significantly outperforms the P control, even in a single-gust environment. Furthermore, we have shown that, in either pivot configuration, an RL control strategy learned in an environment with a finite sequence of gusts can be translated effectively into an environment with an arbitrarily long sequence of gusts.
@@
The present study shows the feasibility of the transformer-based RL, with incorporation pretraining and transfer learning to accelerate training, for control of a highly disturbed flow. However, there are many ways in which the present study can be expanded in future work. For example, we do not investigate the effect of the observation history length, nor the number and locations of the pressure sensors, on the control performance. Additionally, other types of control input (e.g., synthetic jet) and/or other objectives (e.g., drag reduction, non-zero target lift) can be investigated by following the same methodology. These studies could address the underlying question of controllability, which we have not explored in this work.

### v3

@@ III.1 Benchmarking control performance: mid-chord pitching in single-gust environment
We begin by investigating a baseline environment, which is characterized by a single upstream gust. This configuration is set up to be a benchmark for comparing the performance of P control against RL control. Specifically, we intend to evaluate how well each approach can mitigate the lift variation induced by the gust, and to establish a performance baseline before we move on to a more complex environment. The P control strategy, also explored in previous work as a baseline for RL [Beckers2024], is defined as follows:
@@
In principle, the proportional controller used in this study could be extended to a proportional-integral-derivative (PID) controller, and its gains could be tuned, either automatically via some optimization algorithms or manually, against the same reward function used in our RL setting, but each evaluation of a single set of PID gains requires many simulations across different gust scenarios (varying gust parameters and numbers of gust), and a tuning process of PID gains requires many such evaluations, so a systematic optimization of PID gains for all cases considered in this study entails a substantial additional computational effort. Since the focus of this study is not about an optimal PID controller design but rather to compare a simple but well-performing baseline with our RL controller, we restrict attention to a tuned proportional controller and leave a more exhaustive study about PID (and other linear) baselines for future work.
@@ III.2 Extending control capabilities: mid-chord pitching in multi-gust environment
In Fig. 10 we evaluate the trained RL control strategies against the P control strategies over $45$ random cases with three gusts. First, we observe that the case with no control yields a total reward of around 175-175, which is significantly lower than the corresponding value of approximately $15$ in a single-gust environment and reflects the stronger disturbance introduced by three successive gust. Furthermore, the total reward increases with the $K_{p}$ value until it reaches the peak total reward of around 14-14 at $K_{p}=70$. Beyond this point, the performance starts to deteriorate, likely due to the system becoming overly reactive to the gust. Meanwhile, the RL agent shows a clear improvement in performance from the initial pretrained policy (blue dashed line) to the final pretrained policy (blue envelope). The final RL control surpasses the best P control by a reward margin of approximately $7$. This performance gap is notably larger than the marginal $0.4$ improvement observed in the single-gust environment. This improvement shows the superiority of the RL controller, compared to the P controller, in a more challenging multi-gust environment. There are at least two plausible reasons for this improvement: first, the ability of the transformer-based RL controller to exploit the temporal information (i.e. the observation history) becomes even more beneficial in such environments; second, the neural networks may provide a richer nonlinear mapping from observations to actions than a proportional mapping, which may also contribute to this observed gain. A more definitive attribution of this improvement would require a dedicated comparison against linear dynamic controllers with comparable temporal dependencies, which we regard as an interesting direction for future work.
@@
In summary, our results confirm that the performance gap between RL and P control widens in the multi-gust environment. Moreover, we observe that both RL agents trained in single- and three-gust environments can be applied to n-successive gusts effectively, and that the RL agent trained in three gusts generalizes slightly better than the one trained on a single gust, but the margin of improvement is moderate due to the actuation magnitude limit. These insights suggest that further performance gains may be achievable with increased control authority, and thus motivate us to examine the effect of pivot point location on control effectiveness in the next sections.
@@ III.3 Revisiting the pivot location: quarter-chord pitching in single-gust environment
We define the first of these, the added mass contribution, using potential flow theory [inviscidbook]:
@@
The small $K_{p}$ controller also leads to expert data with minimal action variation, which, when used as pretraining data, offers little advantage over the default (random) initialization of the RL policy network. Specifically, since we use zero-center initialization (a standard PyTorch initialization [paszke2019pytorch]) for the policy networks, which produces a near-zero output already, and we employ a fixed initial policy standard deviation of $1$, which allows for substantial exploration early in training, the influence of the expert data is diminished during initial policy updates. Therefore, we chose not to use pretraining in this case. To clarify, our decision to omit pretraining in this configuration is not a limitation of the framework, but rather a deliberate design choice driven by the observed ineffectiveness of expert data in this particular setup. As we will demonstrate, even without pretraining, the RL agent is capable of learning an effective control strategy in the quarter-chord configuration. For the readers who may be interested in pursuing pretraining under these conditions, one potential solution is to develop a full PID controller, which would likely be more successful than P control in reducing the overreaction. One could also apply a scaling factor to amplify the action magnitudes from the expert data before using them for pretraining, which would ensure that the pretrained policy deviates more meaningfully from the zero-center initialized baseline. Another approach may be to employ a smaller initial policy standard deviation, which would constrain the exploration range and help maintain the influence of the pretrained expert data. However, these approaches require careful tuning, and are outside the scope of the present study.
@@
Figure 16: The corresponding vorticity contour and surface pressure distribution at the four steps indicated in Fig. 15. Pressure arrows pointing up represent positive $C_{\Delta p}$, and pointing down represent negative $C_{\Delta p}$.
@@
Figure 18: The corresponding vorticity contour and surface pressure distribution at the four steps indicated in Fig. 17. Pressure arrows pointing up represent positive $C_{\Delta p}$, and pointing down represent negative $C_{\Delta p}$.
@@ IV Conclusions
We have found that P control is an effective control strategy in the mid-chord pitching configuration on its own, but that the performance improvement that RL control exhibits over P control widens for multiple-gust scenarios. In the quarter-chord pitching configuration, we have found that the RL control significantly outperforms the P control, even in a single-gust environment. Furthermore, we have shown that, in either pivot configuration, an RL control strategy learned in an environment with a finite sequence of gusts can be translated effectively into an environment with an arbitrarily long sequence of gusts.
@@
The present study shows the feasibility of the transformer-based RL approach, which incorporates pretraining and transfer learning to accelerate training, for the control of a highly disturbed flow. However, there are many ways in which the present study can be expanded in future work. For example, we do not investigate the effect of the observation history length, nor the number and locations of the pressure sensors, on the control performance. Additionally, other types of control input (e.g., synthetic jet) and/or other objectives (e.g., drag reduction, non-zero target lift) can be investigated by following the same methodology. These studies could address the underlying question of controllability, which we have not explored in this work. Furthermore, we do not explore alternative shapes of the reward function, including the relatively weighting of the lift tracking error term and the actuation penalty term, or other forms of the actuation penalization (e.g. a penalty on control effort). A systematic study of these reward-design choices would also be an interesting direction for future work. Finally, a comprehensive study of the transformer architecture (e.g. different numbers of encoder layers and attention heads) under matched conditions is another promising avenue for future work.
@@ Acknowledgements.
Appendix A A comparison of different observation window sizes
The choice of $N=20$ corresponds to $0.6$ convective time units, which we selected based on the flow physics in this problem setting: we found that the dominant lift variation comes from the vortex-foil interactions near the leading edge, and the boundary layer relaxes relatively quick at $Re=200$ when the gust is passing the foil. Therefore, we expect that $N=20$ already spans the main memory of the disturbance. Furthermore, this window size leads to a relatively lightweight training, which was also empirically demonstrated to enable the RL agent to learn an effective control policy across all gust scenarios considered in the present work, as we discussed in the main text.
To provide a direct comparison, we trained the RL agents with the observation window sizes $N=20$ and $N=40$ in a three-gust environment, as shown in figure 22.
[…]
Figure 22: A comparison of the learning curves for the RL agents having observation window size $N=20$ and $N=40$ in a three-gust environment. The solid lines and the shadow regions are the mean and std values of the episode sum rewards across three different runs. For the dashed lines, the moving average has a window size of $50$. At each episode index, the episode reward is averaged across all parallel environments per run.
We note that the training settings are identical for the two cases (i.e. starting from the same $K_{p}=30$ pretrained policy, the same PPO parameters, the same training hyper-parameters, etc.). As we can observe, the $N=40$ case actually converges slightly slower than the $N=20$, and the mean episode reward for $N=40$ until $150\text{th}$ episode is also slightly lower than the $N=20$. In other words, simply doubling the observation window size without retuning the RL hyper-parameters does not necessarily improve the performance, and can even degrade it. This also corresponds to the above interpretation: once the observation window size covers most of the relevant (useful) information, adding old, weakly informative observations can possibly dilute the learning signal and thus be even harmful for the training without any retuning. However, we do not exclude the possibility that a retuned $N=40$ configuration (e.g. different learning rate, batch size, and the transformer architecture depth and width) may improve upon $N=20$, but given a higher training cost and lack of clear benefit from this experiment, we believe that the choice of $N=20$ is an efficient and sufficient choice in the present study.